OpenStudy (anonymous):

What is the value of x in the equation 3(2x + 8) = 0? –8 –4 4 8

pooja195 (pooja195):

Distribute \[\huge~\rm~(3)(2x)+(3)(8)=0\]

OpenStudy (anonymous):

I don't know how to do that

pooja195 (pooja195):

Whats 3 times 2x ?

pooja195 (pooja195):

Or ignore the x for now whats 3 times 2 ?

OpenStudy (anonymous):

6

pooja195 (pooja195):

Good so we have 6x 3 times 8 =?

OpenStudy (anonymous):

24

pooja195 (pooja195):

6x+24=0 we subtract 24 from both sides so that we can isolate x \[\huge~\rm~~6x+24−24=0−24\] We get our new equation \[\huge~\rm~6x=-24\] 6 div 6 =1 x i alone \[\huge~\rm~-24\div6=?\]

OpenStudy (anonymous):

-4

pooja195 (pooja195):

good so x=-4